OpenStudy (anonymous):

Related rates with an angle: A particle on the x axis is moving to the right at 2 units per second. A second particle on the y axis is moving down at 3 units per second. At a certain instant, the first particle is at the point (5,0) and the second particle is at the point (0,7). How rapidly is the angle between the x axis and the line joining the two particles changing at that instant? Are the particles moving towards or away from each other at that instant?

OpenStudy (anonymous):

let Q be the angle, so, tan(Q) = y/x

OpenStudy (anonymous):

taking the derivative, sec^2(Q) = ((dy/dt)X - (dx/dt)Y) / x^2

OpenStudy (anonymous):

how to solve from there? I get to the point of sec^2(Q) = 5/4

OpenStudy (anonymous):

actually, not sure this is right

OpenStudy (anonymous):

made a mistake, but came up with similar sec^2(Q) = 29/25

Miracrown (miracrown):